# RePo: Language Models with Context Re-Positioning

### v1

@@ 4.2 Results
For all evaluation subtasks, the input contexts are injected with irrelevant information. As shown in Table 1, even when all test contexts are within the training context length (i.e., 4K tokens), RePo outperforms RoPE by 11.04 points. This suggests that the re-positioning mechanism in RePo effectively reduces extraneous cognitive load and improves robustness against noisy or distracting information in model inputs.
@@
The advantage of RePo becomes even more pronounced on long-context tasks. As shown in Figure 2, RePo already surpasses all baselines at a context length of 4K. The performance gap further widens at 8K and 16K, which are lengths unseen during training. To rule out potential confounding effects from RePo’s noise robustness, we additionally evaluate on LongBench (bai-etal-2024-longbench), which consists of more realistic long-context tasks. As shown in Table 3, RePo consistently outperforms other baselines on LongBench by at least 5.48 points. In addition, the hybrid R2N1 method, with interleaved RoPE and NoPE layers, achieves the second-best performance, consistent with the findings of yang2025rope.
Structured Data
Since linearizing structured data (e.g., tables and graphs) into natural language can result in substantial loss of structural information, it is of interest to examine whether contextual re-positioning benefits such data types. As shown in Table 2, RePo improves over the vanilla RoPE method by an average of 1.94 EM points. Interestingly, the NoPE method achieves the best performance on the graph dataset, suggesting that emphasizing local positional relationships may not be a valid assumption for graph-structured inputs.
@@ 5.1 Attention Mass on Relevant Tokens
Since our RePo method re-organizes the context based on its intrinsic structure, we hypothesize that it can better capture long-distance dependencies by bringing distant but relevant tokens with closer positions. To evaluate this effect, we analyze the attention patterns of methods with different types of position assignment strategies on the needle-in-a-haystack (NIAH) task (kamradt2023needle; hsiehruler) and quantitatively measure the attention mass, i.e., attention scores averaged across attention heads and layers, from generated tokens to three non-overlapping parts of the context, following yang2025rope:
@@
We conduct our analysis on the NIAH dataset provided by RULER (hsiehruler), where the context follows the format:
@@ 5.2 Position Patterns Learned by RePo
Table 5: We use the default evaluation suites, including few-shot examples, prompts, and metrics, provided in allenai/olmes for evaluation. All evaluations are conducted within the training context length, i.e., 4096 tokens. We use bold and underline to indicate the best and second-best results, respectively.
Model ARC-C ARC-E BoolQ CoQA Drop Hellaswag MMLU-Pro TriviaQA AVG.
RoPE 47.99 75.25 72.12 56.87 37.90 70.68 13.77 54.98 53.70
NoPE 44.05 73.64 66.70 44.52 33.22 65.68 10.70 43.43 47.74
R2N1 47.30 75.68 73.04 59.31 38.48 69.26 13.41 54.61 53.88
N2R1 43.78 72.72 69.38 50.74 36.66 67.58 11.99 47.31 50.02
RePo 47.61 74.87 73.58 57.44 38.17 70.08 13.52 54.56 53.73
We then analyze the patterns of the positions assigned by RePo. We split the positions $\bm{z}^{l,h}$ into non-overlapping chunks with $\Delta$ tokens $\{\bm{z}^{k,h}_{1:\Delta},\bm{z}^{k,h}_{\Delta+1:2\Delta},\dots,\bm{z}^{k,h}_{L-\Delta:L}\}$ and define three pattern types:
@@
We empirically set $\Delta=16$ and $\epsilon=0.2$ to provide insights for the learned patterns [^5] Setting different values for $\Delta$ and $\epsilon$ results in different plots, but the overall conclusion still holds.. As shown in Figure 4, we find that the Mono pattern is very rare (4% of all chunks), and the model prefer the constant patterns (22% of all chunks) than mono pattern. The dominating pattern of positions assigned by RePo is Hybrid, indicating that the position patterns beneficial for LLMs are different from those pre-defined in previous work (vaswani2017attention; su2024roformer). This analysis is conducted on the RULER benchmark, but we find consistent observations on other benchmarks.
@@ 5.3 Performance on General Tasks
Table 6: Efficiency comparison. FLOPs are reported for training on 50B tokens. Inference time (second) per token is evaluated using vLLM.
@@ 5.4 Efficiency
As shown in Table 6, we compare the FLOPs and inference time between the vanilla model with RoPE method and RePo. We observe that RePo method is very lightweight, introducing only a $0.9\%$ increase in parameters while providing performance gains across many evaluation dimensions. When running inference for RULER benchmark (hsiehruler) within training context length, the time cost of RePo is comparable to that of the vanilla model.
@@ 6 Related Work
The self-attention mechanism in Transformers (vaswani2017attention) is inherently permutation-invariant, lacking an intrinsic understanding of input token order. To address the issue, a position encoding module is used in most Transformer-based models to map the position indices of input tokens into biases or embeddings that can be integrated into the model.
Bias-based position encoding methods, such as ALiBi (press2021train), KERPLE (NEURIPS2022_37a41384), and T5 (raffel2020exploring), integrate a distance bias directly into the attention logits to control the attention field of perception. Most position encoding methods, however, learn embeddings for position indices or distances. The absolute position embedding, which is added to the hidden states, has been widely used in Transformer-based models with small sequence lengths (vaswani2017attention; radford2019language; devlin2019bert). In the era of large language models (LLMs), the RoPE method (su2024roformer), which rotates queries and keys for calculating attention scores based on position distance, has become the de facto choice for modern LLM architectures. Many subsequent variants have been proposed to improve the context extrapolation performance of RoPE (peng2024yarn; chen2023extending; chen2024clex; chen-etal-2025-hope) or manually combining it with NoPE (barbero2025round; yang2025rope). Our work is orthogonal to these enhanced RoPE methods. Recently, li2025seqpe proposed an alternative approach to model positional information in a sequential way.
In contrast to previous literature, our work focuses on position prediction for tokens within a given context, prior to the position encoding process. The most relevant work to ours is COPE (golovneva2024contextual), which learns an attention-logits-based gate module to predict a non-decreasing sequence for positions, aiming to segment high-level abstractions, such as sentences. However, The COPE method uses attention logits for gating, which incurs high time and memory costs due to the computation of $[B,L,L]$ tensors. It is also incompatible with other position encoding methods, like RoPE, and flash attention (dao2024flashattention), limiting its scalability. In contrast, our work focuses on re-organizing tokens within the context using a lightweight RePo module that is compatible with most position encoding methods. Recently, zhang2025agentic also highlighted the importance of input context and proposed ACE, which iteratively evolves the context in an agentic manner. Our work aims to enhance Transformer models with context re-positioning, which is orthogonal to ACE (zhang2025agentic).
@@ 7 Conclusion
In this paper, we addressed high extraneous cognitive load by substituting the rigid linear position encoding in LLMs with context Re-Positioning (RePo). The proposed RePo is a novel, lightweight differentiable module ($f_{\phi}$) that empowers models to learn positions that capture the structure and dependencies in a context. Through continual pre-training of an OLMo-2 1B model, our extensive experiments demonstrated that RePo significantly outperforms strong baselines on tasks requiring specific contextual dependencies, showing substantial gains on noisy-context, structured data, and long-context extrapolation benchmarks, while maintaining performance on general short-context tasks. Our analysis revealed that RePo also brings notable technical advantages, as it mitigates locality bias of traditional position assignment strategies by allocating more attention to distant but relevant ``needle'' tokens, and learns adaptive position patterns in a more non-linear and denser continues space. By enabling models to actively re-position their input context, RePo opens a new direction for flexible context management driven by innovations in LLM architecture.
8 Impact Statement
@@ 9 Acknowledgment
This work was conducted during Huayang's internship at Sakana AI. Huayang wishes to thank his mentors, Richard and Tianyu, for their kind support during this journey. The authors also thank Yoav Gelberg and Yingtao Tian for insightful discussions regarding the experiments, as well as Yujin Tang, Qi Sun, Makoto Shing, Lemao Liu, Deng Cai, Leyang Cui, Yahui Liu, and Tian Lan for their feedback on the manuscript.
@@ Appendix A What’s the Real Difference between Conventional PEs, NoPE, and RePo?
Recently, researchers have found that the causal mask in the attention mechanism enables LLMs to implicitly learn positional information, and that removing explicit positional encoding can even achieve superior performance on structured data and long-context tasks. This approach is referred to as the NoPE method (kazemnejad2023impact; yang2025rope; wang-etal-2024-length; barbero2025round). We argue that the attention score of NoPE can be reformulated within the RoPE framework by assigning a constant positional value $a$:
@@
In addition, LLMs with interpolated NoPE and RoPE layers (yang2025rope; meta2025llama; barbero2025round) have been widely used architectures, which can be explained as hybrid position assignment strategies. In contrast to previous works that empirically configure the hybrid system with hyper-parameters, our RePo shows higher expressiveness, as it can dynamically determine whether to adopt the conventional linear, NoPE-like constant, or hybrid position assignment for tokens in a given context. A comparison among the three approaches is summarized in Table 7. As explained in §2, when $z_{j}=z_{i}$, RePo effectively reduces to the NoPE pattern with the constant $z_{j}=z_{i}=a$. In contrast, if $z_{j}>z_{i}$ for $j>i$, it indicates that RePo adopts positional relationship similar to the conventional linear style, e.g., the strategt used in RoPE. In our experiments and analyses, we will demonstrate that an LLM may dynamically select between constant and linear position assignments, or hybridize them with RePo module. Notably, although we use RoPE for the comparison, linear position assignment is widely adopted in conventional positional encoding methods (vaswani2017attention; gehring2017convolutional; press2021train; li2025seqpe), and our findings can be readily extended to these approaches.
@@ Appendix B Details of Experiments
B.1 Extrapolation
@@ B.2 General Tasks
B.3 Ablation Study
As shown in Figure 5, we evaluate the sensitivity of model performance to the starting layer of RePo, where $l=5$ indicates that RePo is applied beginning from the 5th layer of the LLM, while the vanilla RoPE is used for the lower layers. We conduct experiments on two subtasks, NIAH and MMLUPro. The results show that overall performance is robust to this hyperparameter. However, increasing $l$ slightly improves performance on general benchmarks while negatively affecting performance on NIAH. The results are consistent on other evaluation benchmarks.
@@ C.1 Setup
We use the data and train/dev/test splits provided in kazemnejad2023impact for the text reversal task. The sequence lengths $L$ in training are between $[2,20]$, while we use the sequence lengths between $[2,30]$ for evaluation. The input sequence $\bm{x}$ is constructed from a fixed set of subwords that are shared across the three datasets, without regard to grammatical or semantic structure (kazemnejad2023impact).
We train a GPT-2 model with 4 layers for this task. We use NoPE, RoPE, and RePo  [^6] Notably, RePo functions solely as a position prediction module. For brevity, however, mentioning RePo implies the use of RoPE encoding together in this work. methods to train the model. For the RePo method, we shared the parameters of $f_{\phi}$ for all the attention heads in each layer. All training hyper-parameters are set as in kazemnejad2023impact.
@@ Appendix D Case Study
As shown in Figure 8, we visualize the positions assigned by RePo when testing on the MMLUPro benchmark (wang2024mmlu) with few-shot examples. We observe that RePo learns distinct patterns across different layers and attention heads. Interestingly, as shown in Figure 8(b), the patterns of assigned positions roughly align with the semantic segmentation of the few-shot examples, demonstrating that RePo is capable of capturing the structure of the input context. Additionally, we find that some positions assigned by RePo are negative values, as shown in Figure 8(c). Those negative positions can be interpreted as rotations in a reversed direction under the RoPE framework. There also exist some outlier positions in the figures. Upon inspection, we find that they correspond to non-informative punctuation marks and function words, such as ``.'' and ``such.''

### v2

@@ 4.2 Results
All evaluation subtasks inject substantial irrelevant information into the input context. As shown in Table 1, RePo consistently improves performance under noisy conditions. On OLMo-2 1B, RePo achieves the best average accuracy (91.3), outperforming RoPE by 5.4 points and surpassing alternative re-positioning variants such as R2N1 and N2R1. When scaling to the 7B model, overall performance is already high across all methods; nevertheless, RePo still yields consistent improvements, increasing the average score by 0.6 points over RoPE and achieving perfect accuracy on the multi-value task. These results indicate that the re-positioning mechanism in RePo enhances robustness to noisy context, even when sequence lengths remain within the training regime.
Structured Data
Since linearizing structured data (e.g., tables) into natural language can result in substantial loss of structural information, it is of interest to examine whether contextual re-positioning benefits special data type. As shown in Table 2, RePo substantially improves exact-match accuracy on HybridQA compared to vanilla RoPE. On the 1B model, RePo outperforms RoPE by 2.27 points and exceeding all alternative baselines. The gains are even larger at the 7B scale, where RePo improves over RoPE by 4.09 points. The observations suggest that RePo better preserves and exploits latent structural cues in linearized table representations.
@@
The advantage of RePo becomes even more pronounced on long-context tasks. As shown in Figure 1, RePo based on OLMo-2 1B already surpasses all baselines at a context length of 4K. The performance gap further widens at 8K and 16K, which are lengths unseen during training. To rule out potential confounding effects from RePo’s noise robustness, we additionally evaluate on LongBench (Bai et al., 2024), which consists of more realistic long-context tasks. As reported in Table 3, RePo consistently outperforms RoPE across all task categories. On OLMo-2 1B, RePo improves the average LongBench score by 6.93 points over RoPE. Similar improvements are observed at the 7B scale, where RePo yields a 6.38-point average gain, demonstrating that the proposed re-positioning mechanism scales effectively with model capacity and supports robust generalization to extended contexts.
@@ 5.1 Attention Mass on Relevant Tokens
Since our RePo method re-organizes the context based on its intrinsic structure, we hypothesize that it can better capture long-distance dependencies by bringing distant but relevant tokens with closer positions. To evaluate this effect, we analyze the attention patterns of methods with different types of position assignment strategies on the needle-in-a-haystack (NIAH) task (Kamradt, 2023; Hsieh et al., ) and quantitatively measure the attention mass, i.e., attention scores averaged across attention heads and layers, from generated tokens to three non-overlapping parts of the context, following Yang et al. (2025b):
@@
We conduct our analysis on the NIAH dataset provided by RULER (Hsieh et al., ), where the context follows the format:
@@ 5.2 Position Patterns Learned by RePo
Table 5: We use the default evaluation suites, including few-shot examples, prompts, and metrics, provided in allenai/olmes for evaluation. All evaluations are conducted within the training context length, i.e., 4096 tokens.
Base Method ARC-C ARC-E BoolQ CoQA Drop Hellaswag MMLU-Pro TriviaQA AVG. $\Delta$ vs RoPE
1B RoPE 47.99 75.25 72.12 56.87 37.90 70.68 13.77 54.98 53.70 0.00
NoPE 44.05 73.64 66.70 44.52 33.22 65.68 10.70 43.43 47.74 -5.96
R2N1 47.30 75.68 73.04 59.31 38.48 69.26 13.41 54.61 53.88 ++0.18
N2R1 43.78 72.72 69.38 50.74 36.66 67.58 11.99 47.31 50.02 -3.68
RePo 47.61 74.87 73.58 57.44 38.17 70.08 13.52 54.56 53.73 ++0.03
7B RoPE 78.92 85.74 84.02 69.73 61.07 82.36 27.97 77.56 70.92 0.00
RePo 78.46 85.61 84.32 68.59 59.59 81.92 27.15 76.80 70.30 -0.62
We then analyze the patterns of the positions assigned by RePo. We split the positions $\bm{z}^{k,h}$ into non-overlapping chunks with $\Delta$ tokens $\{\bm{z}^{k,h}_{1:\Delta},\bm{z}^{k,h}_{\Delta+1:2\Delta},\dots,\bm{z}^{k,h}_{L-\Delta:L}\}$ and define three pattern types:
@@
We empirically set $\Delta=16$ and $\epsilon=0.2$ to provide insights for the learned patterns [^5] Setting different values for $\Delta$ and $\epsilon$ results in different plots, but the overall conclusion still holds.. As shown in Figure 3, we find that the Mono pattern is very rare (4% of all chunks), and the model prefer the constant patterns (22% of all chunks) than mono pattern. The dominating pattern of positions assigned by RePo is Hybrid, indicating that the position patterns beneficial for LLMs are different from those pre-defined in previous work (Vaswani et al., 2017; Su et al., 2024). This analysis is conducted on the RULER benchmark, but we find consistent observations on other benchmarks.
@@ 5.3 Performance on General Tasks
Table 6: Efficiency comparison. FLOPs are reported for training OLMo-2 1B model on 50B tokens. Inference time (second) per token is evaluated using vLLM.
@@ 5.4 Efficiency
As shown in Table 6, we compare the FLOPs and inference time between the vanilla model with RoPE method and RePo. We observe that RePo method is very lightweight, introducing only a $0.9\%$ increase in parameters while providing performance gains across many evaluation dimensions. When running inference for RULER benchmark (Hsieh et al., ) within training context length, the time cost of RePo is comparable to that of the vanilla model.
@@ 6 Related Work
The self-attention mechanism in Transformers (Vaswani et al., 2017) is inherently permutation-invariant, lacking an intrinsic understanding of input token order. To address the issue, a position encoding module is used in most Transformer-based models to map the position indices of input tokens into biases or embeddings that can be integrated into the model.
Bias-based position encoding methods, such as ALiBi (Press et al., 2021), KERPLE (Chi et al., 2022), and T5 (Raffel et al., 2020), integrate a distance bias directly into the attention logits to control the attention field of perception. Most position encoding methods, however, learn embeddings for position indices or distances. The absolute position embedding, which is added to the hidden states, has been widely used in Transformer-based models with small sequence lengths (Vaswani et al., 2017; Radford et al., 2019; Devlin et al., 2019). In the era of large language models (LLMs), the RoPE method (Su et al., 2024), which rotates queries and keys for calculating attention scores based on position distance, has become the de facto choice for modern LLM architectures. Many subsequent variants have been proposed to understand and improve the context extrapolation performance of RoPE (Peng et al., 2024; Chen et al., 2023, 2024, 2025; Oka et al., 2025, 2026b, 2026a), and exploring the combination with NoPE (Barbero et al., 2025; Yang et al., 2025b; Gelberg et al., 2025) at different levels. Our work is orthogonal to these enhanced RoPE methods.
In contrast to previous literature, our work focuses on position prediction for tokens within a given context, prior to the position encoding process. The most relevant work to ours is COPE (Golovneva et al., 2024), which learns an attention-logits-based gate module to predict a non-decreasing sequence for positions, aiming to segment high-level abstractions, such as sentences. However, The COPE method uses attention logits for gating, which incurs high time and memory costs due to the computation of $[B,L,L]$ tensors. It is also incompatible with other position encoding methods, like RoPE, and flash attention (Dao, 2024), limiting its scalability. More data-dependent position encoding methods are further designed from the perspective of linear Transformers (Lin et al., 2025; Yang et al., 2025c; Movahedi et al., 2025). In contrast, our work focuses on re-organizing tokens within a context using the lightweight RePo that is compatible with most positional encoding methods and can be applied to standard pre-trained LLMs without training from scratch.
There also exist works that re-organize the context in a model-agnostic way. (Peysakhovich and Lerer, 2023) demonstrate that manually putting the documents with highest attention scores to the end of context achieves more robust performance on long-context question answering. Recently, Zhang et al. (2025) also highlighted the importance of input context and proposed ACE, which iteratively evolves the context in an agentic manner. Our work aims to enhance Transformer architecture with context re-positioning, which is orthogonal to this line.
@@ 7 Conclusion
In this paper, we addressed high extraneous cognitive load by substituting the rigid linear position encoding in LLMs with context Re-Positioning (RePo). The proposed RePo is a novel, lightweight differentiable module ($f_{\phi}$) that empowers models to learn positions that capture the structure and dependencies in a context. Through continual pre-training of an OLMo-2 models, our extensive experiments demonstrated that RePo significantly outperforms strong baselines on tasks requiring specific contextual dependencies, showing substantial gains on noisy-context, structured data, and long-context extrapolation benchmarks, while maintaining performance on general short-context tasks. Our analysis revealed that RePo also brings notable technical advantages, as it mitigates locality bias of traditional position assignment strategies by allocating more attention to distant but relevant ``needle'' tokens, and learns adaptive position patterns in a more non-linear and denser continues space. By enabling models to actively re-position their input context, RePo opens a new direction for flexible context management driven by innovations in LLM architecture.
Impact Statement
@@ 8 Acknowledgment
This work was conducted during Huayang's internship at Sakana AI. Huayang wishes to thank his mentors, Richard and Tianyu, for their kind support during this journey. The authors also thank Yoav Gelberg and Yingtao Tian for insightful discussions regarding the experiments, as well as Yujin Tang, Qi Sun, Makoto Shing, Lemao Liu, Deng Cai, Leyang Cui, Yahui Liu, Tian Lan, and Minghao Wu for their feedback on the manuscript.
@@ Appendix A What’s the Real Difference between Conventional PEs, NoPE, and RePo?
Recently, researchers have found that the causal mask in the attention mechanism enables LLMs to implicitly learn positional information, and that removing explicit positional encoding can even achieve superior performance on structured data and long-context tasks. This approach is referred to as the NoPE method (Kazemnejad et al., 2023; Yang et al., 2025b; Wang et al., 2024a; Barbero et al., 2025). We argue that the attention score of NoPE can be reformulated within the RoPE framework by assigning a constant positional value $a$:
@@
In addition, LLMs with interpolated NoPE and RoPE layers (Yang et al., 2025b; Meta, 2025; Barbero et al., 2025) have been widely used architectures, which can be explained as hybrid position assignment strategies. In contrast to previous works that empirically configure the hybrid system with hyper-parameters, our RePo shows higher expressiveness, as it can dynamically determine whether to adopt the conventional linear, NoPE-like constant, or hybrid position assignment for tokens in a given context. A comparison among the three approaches is summarized in Table 7. As explained in §2, when $z_{j}=z_{i}$, RePo effectively reduces to the NoPE pattern with the constant $z_{j}=z_{i}=a$. In contrast, if $z_{j}>z_{i}$ for $j>i$, it indicates that RePo adopts positional relationship similar to the conventional linear style, e.g., the strategy used in RoPE. In our experiments and analyses, we will demonstrate that an LLM may dynamically select between constant and linear position assignments, or hybridize them with RePo module. Notably, although we use RoPE for the comparison, linear position assignment is widely adopted in conventional positional encoding methods (Vaswani et al., 2017; Gehring et al., 2017; Press et al., 2021; Li et al., 2025), and our findings can be readily extended to these approaches.
@@ Appendix B Details of Experiments
B.1 Noisy Context
We use the following task ids in olmes for the evaluation in Table 1: ruler_niah_s_1__4096::std, ruler_niah_mk_1__4096::std, ruler_niah_mv__4096::std, ruler_niah_mq__4096::std
B.2 Extrapolation
@@ B.3 General Tasks
B.4 Additional Results on Long Context
Results for the context extrapolation on OLMo-2 7B model are shown in Figure 5.
Figure 5: Long-context evaluation on OLMo-2 7B model. We use the data in RULER (Hsieh et al., ) benchmark and apply YaRN (Peng et al., 2024) for all RoPE layers to extend the context.
B.5 Ablation Study
As shown in Figure 4, we evaluate the sensitivity of model performance to the starting layer of RePo on OLMo-2 1B model, where $l=5$ indicates that RePo is applied beginning from the 5th layer of the LLM, while the vanilla RoPE is used for the lower layers. We conduct experiments on two subtasks, NIAH and MMLUPro. The results show that overall performance is robust to this hyperparameter. However, increasing $l$ slightly improves performance on general benchmarks while negatively affecting performance on NIAH. The results are consistent on other evaluation benchmarks.
@@ C.1 Setup
We use the data and train/dev/test splits provided in Kazemnejad et al. (2023) for the text reversal task. The sequence lengths $L$ in training are between $[2,20]$, while we use the sequence lengths between $[2,30]$ for evaluation. The input sequence $\bm{x}$ is constructed from a fixed set of subwords that are shared across the three datasets, without regard to grammatical or semantic structure (Kazemnejad et al., 2023).
We train a GPT-2 model with 4 layers for this task. We use NoPE, RoPE, and RePo  [^6] Notably, RePo functions solely as a position prediction module. For brevity, however, mentioning RePo implies the use of RoPE encoding together in this work. methods to train the model. For the RePo method, we shared the parameters of $f_{\phi}$ for all the attention heads in each layer. All training hyper-parameters are set as in Kazemnejad et al. (2023).
@@ Appendix D Case Study
As shown in Figure 8, we visualize the positions assigned by RePo when testing on the MMLUPro benchmark (Wang et al., 2024b) with few-shot examples. We observe that RePo learns distinct patterns across different layers and attention heads. Interestingly, as shown in Figure 8(b), the patterns of assigned positions roughly align with the semantic segmentation of the few-shot examples, demonstrating that RePo is capable of capturing the structure of the input context. Additionally, we find that some positions assigned by RePo are negative values, as shown in Figure 8(c). Those negative positions can be interpreted as rotations in a reversed direction under the RoPE framework. There also exist some outlier positions in the figures. Upon inspection, we find that they correspond to non-informative punctuation marks and function words, such as ``.'' and ``such.''